OpenStudy (anonymous):

Which of the following is the solution set of the given equation? (x - 3) - 2(x + 6) = -5 {-10} {-4} {8}

Parth (parthkohli):

x - 3 - 2x - 12 = -5 x - 2x = 10 -x = 10 x = -10 1st is correct

OpenStudy (anonymous):

Expansion: \[x - 3 - 2x -12 = -5\] \[-x - 15 = -5\] \[-x = 10\] \[\therefore x = -10\]